# Does FLUX Already Know How to Perform Physically Plausible Image Composition?

### v3

@@ Abstract
Image composition aims to seamlessly insert a user-specified object into a new scene, but existing models struggle with complex lighting (e.g., accurate shadows, water reflections) and diverse, high-resolution inputs. Modern text-to-image diffusion models (e.g., SD3.5, FLUX) already encode essential physical and resolution priors, yet lack a framework to unleash them without resorting to latent inversion, which often locks object poses into contextually inappropriate orientations, or brittle attention surgery. We propose SHINE, a training-free framework for Seamless, High-fidelity Insertion with Neutralized Errors. SHINE introduces manifold-steered anchor loss, leveraging pretrained customization adapters (e.g., IP-Adapter) to guide latents for faithful subject representation while preserving background integrity. Degradation-suppression guidance and adaptive background blending are proposed to further eliminate low-quality outputs and visible seams. To address the lack of rigorous benchmarks, we introduce ComplexCompo, featuring diverse resolutions and challenging conditions such as low lighting, strong illumination, intricate shadows, and reflective surfaces. Experiments on ComplexCompo and DreamEditBench show state-of-the-art performance on standard metrics (e.g., DINOv2) and human-aligned scores (e.g., DreamSim, ImageReward, VisionReward). Code and benchmark will be publicly available upon publication.
@@ 1 Introduction
Image composition, which places a user–specified object into a new scene, is a demanding image editing task. Despite the breathtaking progress of multimodal foundation models (e.g., GPT-5 (OpenAI, 2025), Gemini-2.5 (Gemini2.5, 2025), SeedEdit/Doubao (Shi et al., 2024b), and Grok-4 (gro, 2025)), these generic models still struggle with image composition. Typical failures include imprecise object placement, inconsistent lighting, and the subject’s identity drift (see Fig. 2). These limitations indicate that, as of now, massive multimodal pre-training alone has not yet endowed them with sufficient compositional ability for this task. A natural response has been to train specialized models. Yet building large-scale, high-quality, multi-resolution triplet datasets (object, scene, composite) is prohibitively costly. As a result, most composition models are fine-tuned from base models (e.g., FLUX.1-dev (Black Forest Labs, 2024a), FLUX.1-Fill (Black Forest Labs, 2024c), SDXL (Podell et al., 2024)) on synthetic data generated via inpainting or augmentations (Chen et al., 2024c; Yang et al., 2023; Song et al., 2023; Wang et al., 2025; He et al., 2024).
@@
These models, however, face two main limitations (see Fig. 6): (i) Lighting realism. They struggle to achieve natural composition under complex lighting conditions, such as accurate shadow generation or water reflections for the inserted subject. (ii) Resolution rigidity. They are tied to a fixed resolution, necessitating downsampling or cropping when applied to varied, high-resolution background images, which degrades generation quality. Notably, such issues are absent in the base models, implying that the underlying physical priors are present but are not effectively exploited by fine-tuned variants. The degradation largely stems from low-quality synthetic datasets, which inherit flaws from inpainting models that often mis-handle shadows and reflections, producing implausible edits, hallucinated content, or incomplete object removal (Yu et al., 2025b; Winter et al., 2024).
@@
There have been prior training-free attempts to exploit the priors of text-to-image (T2I) models for advancing image composition, but they fall short for two main reasons. (i) Inversion bottlenecks. Most methods (Lu et al., 2023d; Pham et al., 2024; Yan et al., 2025; Li et al., 2024b) depend on accurate image inversion (Song et al., 2020; Lu et al., 2022; Mokady et al., 2023). In practice, inversion constrains the inserted object to the pose of its reference image, often resulting in contextually inappropriate orientations. Moreover, inversion is less effective for classifier-free guidance (CFG) distilled models (e.g., FLUX), where elevated inversion errors degrade identity preservation. (ii) Fragile attention surgery. Many training-free approaches rely on attention manipulation (Lu et al., 2023d; Yan et al., 2025; Li et al., 2024b). While compatible with the joint self-attention in Multimodal Diffusion Transformers (MMDiT) (Peebles & Xie, 2023), these methods inherit the instability and hyperparameter sensitivity (Lu et al., 2023d), limiting their robustness.
@@
To bridge these gaps we present SHINE, a training-free framework for Seamless, High-fidelity Insertion with Neutralized Errors (see Fig. 1). SHINE comprises three innovations: (i) Manifold-Steered Anchor (MSA) loss, which leverages pretrained open-domain customization adapters (e.g., IP-Adapter (Ye et al., 2023b)) to steer noisy latents toward faithfully representing the reference subject while preserving the structural integrity of the background. (ii) Degradation-Suppression Guidance (DSG) that steers sampling away from low-quality distributions. (iii) Adaptive Background Blending (ABB) that eliminates visible seams along mask boundaries.
@@
Existing benchmarks primarily comprise background images with a fixed resolution of $512\times 512$ pixels. To evaluate performance across diverse, high-resolution, and demanding scenarios, we introduce ComplexCompo, a benchmark that includes varied resolutions, both landscape and portrait orientations, and complex conditions such as low lighting, intense illumination, intricate shadows, and water reflections. Extensive experiments on ComplexCompo and DreamEditBench (Li et al., 2023b) demonstrate that SHINE achieves state-of-the-art (SOTA) performance, surpassing baselines on standard metrics (e.g., DINOv2 (Oquab et al., 2023)) and human-aligned metrics (e.g., DreamSim (Fu et al., 2023), ImageReward (Xu et al., 2023), VisionReward (Xu et al., 2024)).
@@ 2 Related Work
This section reviews prior work on image composition. A more comprehensive discussion, covering image composition, general image editing, and subject-driven generation, is offered in Appendix A. Classical image composition splits into sub-tasks (Niu et al., 2021) such as object placement (Azadi et al., 2020; Zhang et al., 2020a), blending (Wu et al., 2019; Zhang et al., 2020b), harmonization (Cao et al., 2023; Lu et al., 2023b), and shadow generation (Hong et al., 2022; Sheng et al., 2021), typically handled by separate models. Diffusion models have shifted the field toward unified frameworks, either training-based or training-free. Training-based approaches fine-tune diffusion models with curated datasets, adding grounding layers, controllability signals, or identity-preserving supervision from image or video sets (Wang et al., 2025; Chen et al., 2024c; Yang et al., 2023; Song et al., 2023; Lu et al., 2023c). However, they often bias model priors and struggle with complex lighting due to the lack of large-scale real-world triplets. Training-free approaches avoid retraining by manipulating inversion and attention during inference, enabling flexible test-time adaptation (Yan et al., 2025; Li et al., 2024b; 2023b; Lu et al., 2023d; Pham et al., 2024). Yet these methods remain fragile: strong injections preserve identity but fix unnatural poses, while weaker ones improve realism at the cost of fidelity, reflecting a core trade-off between identity preservation and natural composition.
@@ 3 Method
Our framework is built on three core components: Manifold-Steered Anchor (MSA) loss, Degradation-Suppression Guidance (DSG), and Adaptive Background Blending (ABB). Importantly, the design is model-agnostic and requires only standard features of modern generative models: MSA loss assumes that the base model supports either personalization finetuning or provides access to a pretrained personalization adapter, DSG uses self-attention maps, and ABB relies on text-image cross-attention. These mild assumptions enable seamless integration into existing pipelines without architectural changes. We present main results with FLUX, while additional experiments on SDXL (Podell et al., 2024), SD3.5 (Esser et al., 2024), and PixArt (Chen et al., 2023) are provided in Appendix E. The complete algorithm is shown in Algorithm 1.
@@ 3.1 Non-Inversion Latent Preparation
To address these limitations, we abandon inversion and instead perform a one-step forward diffusion to obtain the noisy latent. As illustrated in Fig. 3(a), we use a vision-language model (VLM) (Xue et al., 2024; Chen et al., 2024d; Liu et al., 2024) to caption the subject image and leverage this caption, along with an image inpainting model (Li et al., 2024c; Ju et al., 2024; Zhuang et al., 2024; Black Forest Labs, 2024b), to generate the inpainted background image, denoted as $\bm{x}^{\text{init}}$. The noisy latent is encoded in the VAE space as $\bm{z}^{\text{init}}$ and perturbed to timestep $t\leq T$ via one-step forward diffusion, following the flow matching formulation: $\bm{z}_{t}=(1-\sigma_{t})\bm{z}^{\text{init}}+\sigma_{t}\bm{\epsilon}$, where $\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{1})$.
@@ 3.2 Manifold-Steered Anchor Loss
The Manifold-Steered Anchor (MSA) loss is designed to optimize the noisy latent $\bm{z}_{t}$ (from Sec. 3.1) during the denoising process, steering it toward a reference subject while preserving the structural integrity of the original image. The key intuition is to leverage the prior knowledge embedded in pretrained open-domain customization adapters (or alternatively, personalized LoRAs) such as IP-Adapter (Ye et al., 2023b), PuLID (Guo et al., 2024), and InstantCharacter (Tao et al., 2025), to intervene directly in the diffusion trajectory. Specifically, the MSA loss is defined as:
@@
The Jacobian term necessitates backpropagation through the MMDiT, which is computationally expensive. However, this scenario is analogous to Score Distillation Sampling (SDS) (Poole et al., 2022), where research shows that omitting the Jacobian term yields an effective gradient for optimization with diffusion models. Thus, we adopt the same strategy for optimization.
@@ 3.3 Degradation-Suppression Guidance
1:Input: A background latent $\bm{z}^{\text{bg}}$, a inpainted latent $\bm{z}^{\text{init}}$, a subject latent $\bm{z}^{\text{subj}}$, a user mask $\bm{M}^{\text{user}}$.
@@ 3.4 Adaptive Background Blending
Previous methods typically rely on the user-provided mask $\bm{M}^{\text{user}}$ to preserve the background during each denoising step, blending as $\bm{z}_{t}=\bm{M}^{\text{user}}\odot\bm{z}_{t}+(1-\bm{M}^{\text{user}})\odot\bm{z}^{\text{bg}}_{t}$, but this often introduces visible seams along mask boundaries (see the first row of Fig. 5).
@@
Figure 6: Qualitative comparison of our method with multiple baselines in challenging scenarios, drawn from our benchmark dataset. More qualitative comparisons are available in Appendix K.
@@ 4.1 Experimental Setup
Metrics. Previous methods primarily adapt CLIP-I (Radford et al., 2021) and DINOv2 (Oquab et al., 2023) to assess subject identity consistency. However, these features capture high-level semantic information that may not fully align with human perception of finer-grained attributes. Thus, we further incorporate instance retrieval features (IRF) from (Shao & Cui, 2022) and DreamSim (Fu et al., 2023), which better align with human judgments. An analysis of identity consistency metrics is provided in Appendix I. To evaluate overall image quality, we use ImageReward (IR) (Xu et al., 2023) and VisionReward (VR) (Xu et al., 2024), which are fine-grained, multi-dimensional reward models that more accurately reflect human preferences. For background consistency, LPIPS (Zhang et al., 2018) and SSIM (Wang et al., 2004) are employed.
@@
Implementation Details. In our experiment, we used FLUX.1-dev, a 12B-parameter flow matching model, as the base model, combined with InstantCharacter (Tao et al., 2025) as the adapter (Additional results on SDXL, SD3.5, and PixArt are presented in Appendix E). Our approach also supports per-concept LoRA (Hu et al., 2021a), which requires test-time tuning (Ruiz et al., 2023b) but delivers superior identity consistency compared to an open-domain adapter, making it ideal for scenarios demanding precise identity preservation. The denoising process consists of 20 steps, with the inpainted image perturbed to timestep 15 and denoising initiated from that point. Additional details and hyperparameters are provided in Appendix J.
@@
Table 1: Comparison of composition performance across two benchmarks. The best result in each column is highlighted in bold, while the second-best is underlined. Metrics shown in pink are those specifically trained to better align with human preferences. Abbreviations: IRF: Instance Retrieval Features; IR = ImageReward; VR = VisionReward.
Bench Method
Training
-Free
Base
Model
External
Model
@@
Subject Identity Consistency Background Image Quality
Dream-
Edit-
Bench
MADD (He et al., 2024) SD DINO 0.7118 0.6279 0.4333 0.5810 0.0604 0.8182 -0.2545 2.7011
ObjectStitch (Song et al., 2023) SD VIT 0.7567 0.6930 0.5525 0.5093 0.0190 0.8316 0.0791 3.2416
DreamCom (Lu et al., 2023c) SD LoRA 0.7414 0.6749 0.5597 0.5626 0.0200 0.8283 0.1873 3.5053
AnyDoor (Chen et al., 2024c) SD DINO 0.8183 0.7283 0.7714 0.3764 0.0251 0.8894 0.4511 3.3946
UniCombine (Wang et al., 2025) FLUX LoRA 0.8058 0.7332 0.7579 0.3984 0.0050 0.9397 0.4565 3.6108
PBE (Yang et al., 2023) SD - 0.7742 0.7040 0.5845 0.4985 0.0197 0.8287 0.2083 3.3482
TIGIC (Li et al., 2024b) SD - 0.7226 0.6718 0.4711 0.6108 0.0584 0.8153 -0.1332 2.9873
TALE (Pham et al., 2024) SD - 0.7329 0.6604 0.5007 0.6176 0.0392 0.8251 -0.1502 3.1349
TF-ICON (Lu et al., 2023d) SD - 0.7479 0.6865 0.5179 0.5441 0.0582 0.8111 0.0816 3.2823
DreamEdit (Li et al., 2023b) SD LoRA, VIT 0.7703 0.7151 0.6147 0.5047 0.0140 0.9775 0.1744 3.1775
EEdit (Yan et al., 2025) FLUX - 0.6998 0.6590 0.4438 0.6160 0.0039 0.9475 0.0216 3.3606
Ours-Adapter FLUX Adapter 0.8086 0.7415 0.7702 0.3730 0.0236 0.8959 0.5709 3.6234
Ours-LoRA FLUX LoRA 0.8125 0.7452 0.7900 0.3577 0.0271 0.8847 0.5906 3.6161
Complex-
Compo
MADD (He et al., 2024) SD DINO 0.6780 0.5993 0.3638 0.5979 0.0781 0.5658 -0.0088 2.6582
ObjectStitch (Song et al., 2023) SD VIT 0.7608 0.7077 0.5513 0.4717 0.0388 0.6357 0.2482 3.4411
DreamCom (Lu et al., 2023c) SD LoRA 0.648 0.5692 0.2788 0.8192 0.0389 0.6342 -0.0778 3.4409
AnyDoor (Chen et al., 2024c) SD DINO 0.7982 0.7052 0.7319 0.4493 0.0299 0.7262 0.3804 3.3787
UniCombine (Wang et al., 2025) FLUX LoRA 0.7361 0.6552 0.5380 0.5682 0.0237 0.7077 0.2470 3.5454
PBE (Yang et al., 2023) SD - 0.7537 0.6802 0.5189 0.5187 0.0397 0.6321 0.2139 3.4310
TIGIC (Li et al., 2024b) SD - 0.6913 0.6329 0.3848 0.6549 0.0929 0.6228 -0.131 2.8898
TALE (Pham et al., 2024) SD - 0.6816 0.6151 0.3799 0.6773 0.059 0.6334 0.0783 3.4498
TF-ICON (Lu et al., 2023d) SD - 0.6987 0.6435 0.4167 0.6030 0.0815 0.6216 0.1798 3.4323
DreamEdit (Li et al., 2023b) SD LoRA, VIT 0.7314 0.6722 0.5069 0.5670 0.0468 0.7201 0.1212 3.2531
EEdit (Yan et al., 2025) FLUX - 0.6713 0.6153 0.3797 0.6821 0.0226 0.7107 0.1433 3.5009
Ours-Adapter FLUX Adapter 0.7721 0.7107 0.6764 0.4294 0.0404 0.7789 0.4090 3.6020
Ours-LoRA FLUX LoRA 0.7999 0.7384 0.7659 0.3542 0.0430 0.7634 0.4246 3.5951

### v4

# arXiv:2509.21278v4[cs.CV] 28 Feb 2026

@@ Abstract
Image composition aims to seamlessly insert a user-specified object into a new scene, but existing models struggle with complex lighting (e.g., accurate shadows, water reflections) and diverse, high-resolution inputs. Modern text-to-image diffusion models (e.g., SD3.5, FLUX) already encode essential physical and resolution priors, yet lack a framework to unleash them without resorting to latent inversion, which often locks object poses into contextually inappropriate orientations, or brittle attention surgery. We propose SHINE, a training-free framework for Seamless, High-fidelity Insertion with Neutralized Errors. SHINE introduces manifold-steered anchor loss, leveraging pretrained customization adapters (e.g., IP-Adapter) to guide latents for faithful subject representation while preserving background integrity. Degradation-suppression guidance and adaptive background blending are proposed to further eliminate low-quality outputs and visible seams. To address the lack of rigorous benchmarks, we introduce ComplexCompo, featuring diverse resolutions and challenging conditions such as low lighting, strong illumination, intricate shadows, and reflective surfaces. Experiments on ComplexCompo and DreamEditBench show state-of-the-art performance on standard metrics (e.g., DINOv2) and human-aligned scores (e.g., DreamSim, ImageReward, VisionReward). Code is available at https://github.com/ZhumingLian/SHINE.
@@ 1 Introduction
Image composition, which places a user–specified object into a new scene, is a demanding image editing task. Despite the breathtaking progress of multimodal foundation models (e.g., GPT-5 (OpenAI, 2025), Gemini-2.5 (Gemini2.5, 2025), SeedEdit/Doubao (Shi et al., 2024b), and Grok-4 (gro, 2025)), these generic models still struggle with image composition. Typical failures include imprecise object placement, inconsistent lighting, and the subject’s identity drift (see Fig. 2). These limitations indicate that, as of now, massive multimodal pre-training alone has not yet endowed them with sufficient compositional ability for this task. A natural response has been to train specialized models. Yet building large-scale, high-quality, multi-resolution triplet datasets (object, scene, composite) is prohibitively costly. As a result, most composition models are fine-tuned from base models (e.g., FLUX.1-dev (Black Forest Labs, 2024a), FLUX.1-Fill (Black Forest Labs, 2024c), SDXL (Podell et al., 2024)) on synthetic data generated via inpainting or augmentations (Chen et al., 2024c; Yang et al., 2023; Song et al., 2023; Wang et al., 2025a; He et al., 2024).
@@
These models, however, face two main limitations (see Fig. 6): (i) Lighting realism. They struggle to achieve natural composition under complex lighting conditions, such as accurate shadow generation or water reflections for the inserted subject. (ii) Resolution rigidity. They are tied to a fixed resolution, necessitating downsampling or cropping when applied to varied, high-resolution background images, which degrades generation quality. Notably, such issues are absent in the base models, implying that the underlying physical priors are present but are not effectively exploited by fine-tuned variants. The degradation largely stems from low-quality synthetic datasets, which inherit flaws from inpainting models that often mis-handle shadows and reflections, producing implausible edits, hallucinated content, or incomplete object removal (Yu et al., 2025c; Winter et al., 2025).
@@
There have been prior training-free attempts to exploit the priors of text-to-image (T2I) models for advancing image composition, but they fall short for two main reasons. (i) Inversion bottlenecks. Most methods (Lu et al., 2023d; Pham et al., 2024; Yan et al., 2025; Li et al., 2024b) depend on accurate image inversion (Song et al., 2021; Lu et al., 2022; Mokady et al., 2023). In practice, inversion constrains the inserted object to the pose of its reference image, often resulting in contextually inappropriate orientations. Moreover, inversion is less effective for classifier-free guidance (CFG) distilled models (e.g., FLUX), where elevated inversion errors degrade identity preservation. (ii) Fragile attention surgery. Many training-free approaches rely on attention manipulation (Lu et al., 2023d; Yan et al., 2025; Li et al., 2024b). While compatible with the joint self-attention in Multimodal Diffusion Transformers (MMDiT) (Peebles & Xie, 2023), these methods inherit the instability and hyperparameter sensitivity (Lu et al., 2023d), limiting their robustness.
@@
To bridge these gaps we present SHINE, a training-free framework for Seamless, High-fidelity Insertion with Neutralized Errors (see Fig. 1). SHINE comprises three innovations: (i) Manifold-Steered Anchor (MSA) loss, which leverages pretrained open-domain customization adapters (e.g., IP-Adapter (Ye et al., 2023)) to steer noisy latents toward faithfully representing the reference subject while preserving the structural integrity of the background. (ii) Degradation-Suppression Guidance (DSG) that steers sampling away from low-quality distributions. (iii) Adaptive Background Blending (ABB) that eliminates visible seams along mask boundaries.
@@
Existing benchmarks primarily comprise background images with a fixed resolution of $512\times 512$ pixels. To evaluate performance across diverse, high-resolution, and demanding scenarios, we introduce ComplexCompo, a benchmark that includes varied resolutions, both landscape and portrait orientations, and complex conditions such as low lighting, intense illumination, intricate shadows, and water reflections. Extensive experiments on ComplexCompo and DreamEditBench (Li et al., 2023b) demonstrate that SHINE achieves state-of-the-art (SOTA) performance, surpassing baselines on standard metrics (e.g., DINOv2 (Oquab et al., 2024)) and human-aligned metrics (e.g., DreamSim (Fu et al., 2023), ImageReward (Xu et al., 2023), VisionReward (Xu et al., 2024)).
@@ 2 Related Work
This section reviews prior work on image composition. A more comprehensive discussion, covering image composition, general image editing, and subject-driven generation, is offered in Appendix A. Classical image composition splits into sub-tasks (Niu et al., 2021) such as object placement (Azadi et al., 2020; Zhang et al., 2020a), blending (Wu et al., 2019; Zhang et al., 2020b), harmonization (Cao et al., 2023; Lu et al., 2023b), and shadow generation (Hong et al., 2022; Sheng et al., 2021), typically handled by separate models. Diffusion models have shifted the field toward unified frameworks, either training-based or training-free. Training-based approaches fine-tune diffusion models with curated datasets, adding grounding layers, controllability signals, or identity-preserving supervision from image or video sets (Wang et al., 2025a; Chen et al., 2024c; Yang et al., 2023; Song et al., 2023; Lu et al., 2023c). However, they often bias model priors and struggle with complex lighting due to the lack of large-scale real-world triplets. Training-free approaches avoid retraining by manipulating inversion and attention during inference, enabling flexible test-time adaptation (Yan et al., 2025; Li et al., 2024b; 2023b; Lu et al., 2023d; Pham et al., 2024). Yet these methods remain fragile: strong injections preserve identity but fix unnatural poses, while weaker ones improve realism at the cost of fidelity, reflecting a core trade-off between identity preservation and natural composition.
@@ 3 Method
Our framework is built on three core components: Manifold-Steered Anchor (MSA) loss, Degradation-Suppression Guidance (DSG), and Adaptive Background Blending (ABB). Importantly, the design is model-agnostic and requires only standard features of modern generative models: MSA loss assumes that the base model supports either personalization finetuning or provides access to a pretrained personalization adapter, DSG uses self-attention maps, and ABB relies on text-image cross-attention. These mild assumptions enable seamless integration into existing pipelines without architectural changes. We present main results with FLUX, while additional experiments on SDXL (Podell et al., 2024), SD3.5 (Esser et al., 2024), and PixArt (Chen et al., ) are provided in Appendix E. The complete algorithm is shown in Algorithm 1.
@@ 3.1 Non-Inversion Latent Preparation
To address these limitations, we abandon inversion and instead perform a one-step forward diffusion to obtain the noisy latent. As illustrated in Fig. 3(a), we use a vision-language model (VLM) (Xue et al., 2024; Chen et al., 2024d; Liu et al., 2024) to caption the subject image and leverage this caption, along with an image inpainting model (Li et al., 2024c; Ju et al., 2024; Zhuang et al., 2024; Black Forest Labs, 2024b), to generate the image to which the subject is attached, denoted as $\bm{x}^{\text{init}}$. The noisy latent is encoded in the VAE space as $\bm{z}^{\text{init}}$ and perturbed to timestep $t\leq T$ via one-step forward diffusion, following the flow matching formulation: $\bm{z}_{t}=(1-\sigma_{t})\bm{z}^{\text{init}}+\sigma_{t}\bm{\epsilon}$, where $\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{1})$.
@@ 3.2 Manifold-Steered Anchor Loss
The Manifold-Steered Anchor (MSA) loss is designed to optimize the noisy latent $\bm{z}_{t}$ (from Sec. 3.1) during the denoising process, steering it toward a reference subject while preserving the structural integrity of the original image. The key intuition is to leverage the prior knowledge embedded in pretrained open-domain customization adapters (or alternatively, personalized LoRAs) such as IP-Adapter (Ye et al., 2023), PuLID (Guo et al., 2024), and InstantCharacter (Tao et al., 2025), to intervene directly in the diffusion trajectory. Specifically, the MSA loss is defined as:
@@
The Jacobian term necessitates backpropagation through the MMDiT, which is computationally expensive. However, this scenario is analogous to Score Distillation Sampling (SDS) (Poole et al., ), where research shows that omitting the Jacobian term yields an effective gradient for optimization with diffusion models. Thus, we adopt the same strategy for optimization.
@@ 3.3 Degradation-Suppression Guidance
1:Input: A background latent $\bm{z}^{\text{bg}}=\text{VAE}(\bm{x}^{\text{bg}})$, a subject latent $\bm{z}^{\text{subj}}=\text{VAE}(\bm{x}^{\text{subj}})$, a inpainted latent $\bm{z}^{\text{init}}=\text{VAE}\left(\text{Inpainting}\left[\bm{x}^{\text{bg}},\bm{M}^{\text{user}},\text{VLM}(\bm{x}^{\text{subj}})\right]\right)$, a user mask $\bm{M}^{\text{user}}$.
@@ 3.4 Adaptive Background Blending
Previous methods typically rely on the user-provided mask $\bm{M}^{\text{user}}$ to preserve the background during each denoising step, blending as $\bm{z}^{\prime}_{t}=\bm{M}^{\text{user}}\odot\bm{z}_{t}+(1-\bm{M}^{\text{user}})\odot\bm{z}^{\text{bg}}_{t}$, but this often introduces visible seams along mask boundaries (see the first row of Fig. 5).
@@
Figure 6: Qualitative comparison of our method with multiple baselines in challenging scenarios, drawn from our benchmark dataset. More qualitative comparisons are available in Appendix O.
@@ 4.1 Experimental Setup
Metrics. Previous methods primarily adapt CLIP-I (Radford et al., 2021) and DINOv2 (Oquab et al., 2024) to assess subject identity consistency. However, these features capture high-level semantic information that may not fully align with human perception of finer-grained attributes. Thus, we further incorporate instance retrieval features (IRF) from (Shao & Cui, 2022) and DreamSim (Fu et al., 2023), which better align with human judgments. An analysis of identity consistency metrics is provided in Appendix I. For overall image quality, we use ImageReward (IR) (Xu et al., 2023) and VisionReward (VR) (Xu et al., 2024), fine-grained reward models that more accurately reflect human preferences. To more comprehensively evaluate composition quality, we further include three UnifiedReward variants (Wang et al., 2025c; d) and HPSv3 (Ma et al., 2025b) in Appendix L. Background consistency is measured using LPIPS (Zhang et al., 2018) and SSIM (Wang et al., 2004).
@@
Implementation Details. In our experiment, we used FLUX.1-dev, a 12B-parameter flow matching model, as the base model, combined with InstantCharacter (Tao et al., 2025) as the adapter (Additional results on SDXL, SD3.5, and PixArt are presented in Appendix E). Our approach also supports per-concept LoRA (Hu et al., ), which requires test-time tuning (Ruiz et al., 2023) but delivers superior identity consistency compared to an open-domain adapter, making it ideal for scenarios demanding precise identity preservation. The denoising schedule consists of 20 steps, with the inpainted image perturbed to timestep 15 and denoising initiated from that point. We use Flux.1 Fill (Black Forest Labs, 2024b) as the inpainting model and BLIP-3 (Xue et al., 2024) as the VLM. Additional details and hyperparameters are provided in Appendix J.
@@
Table 1: Comparison of composition performance across two benchmarks. The best result in each column is highlighted in bold, while the second-best is underlined. Metrics shown in pink are those specifically trained to better align with human preferences. Abbreviations: IRF= Instance Retrieval Features; IR = ImageReward; VR = VisionReward; URE = UnifiedReward-Edit-qwen3vl-8b.
@@
Bench Method Training -Free Base Model External Model Subject Identity Consistency Background Image Quality
CLIP-I $\uparrow$ DINOv2 $\uparrow$ IRF $\uparrow$ DreamSim $\downarrow$ LPIPS $\downarrow$ SSIM $\uparrow$ IR $\uparrow$ VR $\uparrow$ HPS $\uparrow$ URE $\uparrow$
Dream- Edit- Bench (220) Flux.1 Fill (Black Forest Labs, 2024b) FLUX - 0.7328 0.6745 0.5754 0.5233 0.0166 0.9076 0.5577 3.5997 8.6432 21.5812
MADD (He et al., 2024) SD DINO 0.7118 0.6279 0.4333 0.5810 0.0604 0.8182 -0.2545 2.7011 1.2443 13.8148
ObjectStitch (Song et al., 2023) SD VIT 0.7567 0.6930 0.5525 0.5093 0.0190 0.8316 0.0791 3.2416 7.4529 19.1886
DreamCom (Lu et al., 2023c) SD LoRA 0.7414 0.6749 0.5597 0.5626 0.0200 0.8283 0.1873 3.5053 5.9324 19.9296
AnyDoor (Chen et al., 2024c) SD DINO 0.8183 0.7283 0.7714 0.3764 0.0251 0.8894 0.4511 3.3946 8.4867 19.0989
UniCombine (Wang et al., 2025a) FLUX LoRA 0.8058 0.7332 0.7579 0.3984 0.0050 0.9397 0.4565 3.6108 8.8415 21.7080
PBE (Yang et al., 2023) SD - 0.7742 0.7040 0.5845 0.4985 0.0197 0.8287 0.2083 3.3482 8.3789 20.2137
TIGIC (Li et al., 2024b) SD - 0.7226 0.6718 0.4711 0.6108 0.0584 0.8153 -0.1332 2.9873 5.2676 17.1000
TALE (Pham et al., 2024) SD - 0.7329 0.6604 0.5007 0.6176 0.0392 0.8251 -0.1502 3.1349 6.3773 18.0784
TF-ICON (Lu et al., 2023d) SD - 0.7479 0.6865 0.5179 0.5441 0.0582 0.8111 0.0816 3.2823 7.2643 18.2716
DreamEdit (Li et al., 2023b) SD LoRA, VIT 0.7703 0.7151 0.6147 0.5047 0.0140 0.9775 0.1744 3.1775 6.0250 15.7636
EEdit (Yan et al., 2025) FLUX - 0.6998 0.6590 0.4438 0.6160 0.0039 0.9475 0.0216 3.3606 6.6689 19.5603
Ours-Adapter FLUX Adapter 0.8086 0.7415 0.7702 0.3730 0.0236 0.8959 0.5709 3.6234 8.8861 22.0182
Ours-LoRA FLUX LoRA 0.8125 0.7452 0.7900 0.3577 0.0271 0.8847 0.5906 3.6161 8.8688 21.9421
Complex- Compo (300) Flux.1 Fill (Black Forest Labs, 2024b) FLUX - 0.7108 0.6475 0.5466 0.6018 0.0232 0.7442 0.4088 3.5737 8.7376 19.7712
MADD (He et al., 2024) SD DINO 0.6780 0.5993 0.3638 0.5979 0.0781 0.5658 -0.0088 2.6582 5.9673 13.0567
ObjectStitch (Song et al., 2023) SD VIT 0.7608 0.7077 0.5513 0.4717 0.0388 0.6357 0.2482 3.4411 8.8389 18.8283
DreamCom (Lu et al., 2023c) SD LoRA 0.648 0.5692 0.2788 0.8192 0.0389 0.6342 -0.0778 3.4409 7.9884 18.6143
AnyDoor (Chen et al., 2024c) SD DINO 0.7982 0.7052 0.7319 0.4493 0.0299 0.7262 0.3804 3.3787 8.9760 18.3550
UniCombine (Wang et al., 2025a) FLUX LoRA 0.7361 0.6552 0.5380 0.5682 0.0237 0.7077 0.2470 3.5454 8.8999 19.8529
PBE (Yang et al., 2023) SD - 0.7537 0.6802 0.5189 0.5187 0.0397 0.6321 0.2139 3.4310 8.5923 18.9507
TIGIC (Li et al., 2024b) SD - 0.6913 0.6329 0.3848 0.6549 0.0929 0.6228 -0.131 2.8898 7.6630 16.4301
TALE (Pham et al., 2024) SD - 0.6816 0.6151 0.3799 0.6773 0.059 0.6334 0.0783 3.4498 8.7351 18.7567
TF-ICON (Lu et al., 2023d) SD - 0.6987 0.6435 0.4167 0.6030 0.0815 0.6216 0.1798 3.4323 9.3258 18.2366
DreamEdit (Li et al., 2023b) SD LoRA, VIT 0.7314 0.6722 0.5069 0.5670 0.0468 0.7201 0.1212 3.2531 8.0434 15.8934
EEdit (Yan et al., 2025) FLUX - 0.6713 0.6153 0.3797 0.6821 0.0226 0.7107 0.1433 3.5009 8.7835 19.7348
Ours-Adapter FLUX Adapter 0.7721 0.7107 0.6764 0.4294 0.0404 0.7789 0.4090 3.6020 9.6485 20.7349
Ours-LoRA FLUX LoRA 0.7999 0.7384 0.7659 0.3542 0.0430 0.7634 0.4246 3.5951 9.8418 21.0326